# Trans-dimensional Hamiltonian model selection and parameter estimation from sparse, noisy data

### v1

@@ Appendix C Experimental procedure
CPMG data were measured on a single, naturally occurring NV center in a high-purity chemical-vapor-deposition homoepitaxial diamond (type IIa, 1.1% natural abundance of 13C) with $\langle 111\rangle$ crystal orientation. A solid immersion lens with an aluminum-oxide anti-reflection coating (atomic-layer deposition) was used for enhanced photon collection. Microwave fields for electron spin manipulation were delivered via on-chip lithographically-defined strip lines, and a static magnetic field of $B_{z}\approx 403$ G was applied along the NV-axis using a room-temperature neodymium magnet, with stability $<3$ mG and alignment within $0.07^{\circ}$ (thermal echo calibration). The electron spin Rabi frequency was $14.31(3)$ MHz, with Hermite pulse shapes for efficient MW pulses, and phase alternation of $\pi$-pulses was performed using the XY-8 scheme. Experiments were conducted at $3.7$ K in a commercial closed-cycle cryostat (Montana Cryostation), achieving single-shot NV spin-state readout fidelity of $94.5\%$ via spin-selective resonant excitation. The electron spin relaxation time was $T_{1}>1$ h, natural dephasing time $T_{2}=4.9(2)$ $\mu$s, spin-echo coherence time $T_{2}=1.182(5)$ ms, and multipulse dynamical decoupling coherence time $T_{2,\text{DD}}>1$ s with optimized inter-pulse delay $2\tau$.
@@ Appendix D Dependence on hyperfine couplings
Because our model currently constrains the possible hyperfine couplings to those computed from ab initio density functional theory for each lattice site, it is essential to understand how the recovery method performs when these input couplings are inaccurate. We investigated the hybrid algorithm’s robustness to perturbations in hyperfine couplings, focusing on the ability to recover the hyperfine parameters $A_{\perp}$ and $A_{\parallel}$ from noisy input hyperfine data associated with each lattice site. Simulated inaccuracies in hyperfine couplings, introduced by perturbations $A_{\perp}\pm\delta$ and $A_{\parallel}\pm\delta$, were considered, where we varied the perturbation (denoted as $\delta$) to assess recovery accuracy. We applied our hybrid recovery method to simulations with a 0.001 shot noise per data point, 250 $\tau$ sampled, a 0.008 ms longest $\tau$, 16 CP pulses, and 311 G magnetic field/
Our results indicate that the algorithm remains robust to hyperfine perturbations up to approximately 1 kHz. When the input hyperfine couplings were accurate to the closest kHz, the algorithm performed well, with recovery rates close to 1 for the highest magnitude of hyperfine couplings ($>150$ kHz). However, when the perturbations exceed 1 kHz, the recovery rates for both hyperfine couplings and number of spins decreases, suggesting that the method is less effective when hyperfine couplings are known with low precision (Fig. 9). This threshold suggests that the method is particularly suited for systems where hyperfine couplings can be known to within a kilohertz. Given that state-of-the-art ab initio calculations based on density functional theory can achieve relative mean errors of around 1% for hyperfine couplings at all distances for an NV center in diamond [34], and that we seek to detect nuclear spins in the weakly coupled regime (up to 200 kHz), with the majority of spins considered here (over 99%) below 100 kHz, this level of accuracy falls within the algorithm’s robustness threshold and is promising for practical applications.
@@ Appendix F Detection rates for hyperfine couplings extracted from experimental data
We present the detection rates for hyperfine couplings (Table LABEL:tab:hf_compare) and spatial positions (Table LABEL:tab:positions_compare) of nuclear spins, computed from the posterior distribution sampled using the hybrid MCMC algorithm applied to 250 data points from an $N=32$ CPMG experiment with $\tau$ ranging from 6 to 8 $\mu$s. These results are compared with experimentally reported values in Refs. [16, 35]. We align the computed lattice sites and hyperfine couplings, which were computed in Ref. [34], using the HSE06 functional with a supercell containing 1728 atoms, with the experimentally reported values in Ref. [35]. If an experimental nuclear spin position is reported with low precision, meaning that multiple lattice sites fall within the range defined by the standard deviation, we identify the “recovered" lattice site as the single lattice position within this range with the effective spin frequency ($A_{\text{rec}}^{-}$) that most closely matches the experimentally determined effective spin frequency ($A_{\text{exp}}^{-}$). There are two nuclear spins, C37 and C49, for which we could not find a good match between the experimentally reported positions and effective spin frequencies with the computed positions and effective spin frequencies.
As noted in Ref. [34], there is a known sign inconsistency between the experimentally reported parallel components of hyperfine values and the values computed theoretically. To maintain consistency, we use the sign convention of the experimental data. The hyperfine components computed with the HSE06 hybrid functional in Ref. [34], which we also use in this work, were validated against three separate experimental datasets: Refs. [52, 10, 35]. One of these datasets is the same as the one we use for validation here. Because the HSE06 functional was tested on multiple datasets, we consider it a reliable choice for this study and are not overly concerned about overfitting to a single dataset.
@@
Table 2: Comparison of Experimentally Measured and Recovered Hyperfine Interaction Values. This table compares experimentally measured and recovered hyperfine interaction values for nuclear spins, following the labeling convention in [16, 35]. We report the experimentally measured effective nuclear spin frequencies $A_{\text{exp}}^{-}=\sqrt{(\gamma B_{z}-A_{\parallel,\text{exp}})^{2}+A_{\perp,\text{exp}}^{2}}$ from [35], as well as the parallel ($A_{\parallel,\text{exp}}$) and perpendicular ($A_{\perp,\text{exp}}$) components of the hyperfine interaction for nuclear spins C1–C27 from [16]. These values are compared with the recovered frequencies $A_{\text{rec}}^{-}$ and hyperfine components ($A_{\parallel,\text{rec}},A_{\perp,\text{rec}}$) obtained using a hybrid MCMC algorithm (five ensembles of 25,000 steps each, with the first 10,000 discarded as burn-in), applied to 250 data points from an $N=32$ CPMG experiment with $\tau$ ranging from 6 to 8 $\mu$s. Detection rates of individual nuclear spins are determined based on the frequency of their appearance in spin configurations sampled from the posterior distribution. Spins can only be identified up to a degree of symmetry determined by the number of lattice sites with identical spins; however, we properly account for the multiplicity of spins with identical hyperfine couplings but different lattice sites when computing detection rates. Spins with detection rates above their baseline probabilities are marked with an asterisk (*).

### v2

@@ Appendix C Experimental procedure
[…]
@@ Appendix D Dependence on hyperfine couplings
Because our model currently constrains the possible hyperfine couplings to those computed from ab initio density functional theory for each lattice site, it is essential to understand how the recovery method performs when these input couplings are inaccurate. We investigated the hybrid algorithm’s robustness to perturbations in hyperfine couplings, focusing on the ability to recover the hyperfine parameters $A_{\perp}$ and $A_{\parallel}$ from noisy input hyperfine data associated with each lattice site. Simulated inaccuracies in hyperfine couplings, introduced by perturbations $A_{\perp}\pm\delta$ and $A_{\parallel}\pm\delta$, were considered, where we varied the perturbation (denoted as $\delta$) to assess recovery accuracy. We applied our hybrid recovery method to simulations with a 0.001 shot noise per data point, 250 $\tau$ sampled, a 0.008 ms longest $\tau$, 16 CP pulses, and 311 G magnetic field.
Our results indicate that the algorithm remains robust to hyperfine perturbations up to approximately 1 kHz. When the input hyperfine couplings were accurate to the closest kHz, the algorithm performed well, with recovery rates close to 1 for the highest magnitude of hyperfine couplings (>150>150 kHz). However, when the perturbations exceed 1 kHz, the recovery rates for both hyperfine couplings and number of spins decreases, suggesting that the method is less effective when hyperfine couplings are known with low precision (Fig. 9). This threshold suggests that the method is particularly suited for systems where hyperfine couplings can be known to within a kilohertz. Given that state-of-the-art ab initio calculations based on density functional theory can achieve relative mean errors of around 1% for hyperfine couplings at all distances for an NV center in diamond [34], and that we seek to detect nuclear spins in the weakly coupled regime (up to 200 kHz), with the majority of spins considered here (over 99%) below 100 kHz, this level of accuracy falls within the algorithm’s robustness threshold and is promising for practical applications.
@@ Appendix F Detection rates for hyperfine couplings extracted from experimental data
[…]
As noted in Ref. [34], there is a known sign inconsistency between the experimentally reported parallel components of hyperfine values and the values computed theoretically. To maintain consistency, we use the sign convention of the experimental data. The hyperfine components computed with the HSE06 hybrid functional in Ref. [34], which we also use in this work, were validated against three separate experimental datasets: Refs. [56, 10, 35]. One of these datasets is the same as the one we use for validation here. Because the HSE06 functional was tested on multiple datasets, we consider it a reliable choice for this study and are not overly concerned about overfitting to a single dataset. In Fig. 5(d), the yellow points correspond to all posterior hyperfines sampled by the MCMC algorithm, whereas the crosses indicate experimentally observed spins. The crosses plotted in Fig. 5(d) correspond to columns 7 and 8 (rather than columns 4 and 5) in Table LABEL:tab:hf_compare, which was chosen to better visualize the correspondence between spins that could be recovered and the posterior hyperfine distribution. That is, all plotted crosses correspond to DFT-computed hyperfines, and in Tables 2 and 3 we provide the subset of spins for which we consider the DFT-computed hyperfines and experimentally measured hyperfines to be equivalent. The matching criterion uses both lattice positions and hyperfine couplings to identify corresponding spins.
We note that the 50 spins associated with this NV center were characterized using different types of experimental measurements, and therefore different amounts of hyperfine information are available across the set. All spins have reconstructed lattice coordinates (Table 3), which we use as the primary criterion to associate experimentally observed spins with DFT lattice sites for hyperfine comparison. For spins 1–27, both $m_{s}=\pm 1$ transition frequencies were measured, allowing direct extraction of $A_{\parallel}$ and $A_{\perp}$ (Table 2). For more information on the experimental procedures for spins 1-27, see [16]. For spins 28–50, only partial spectroscopic data are available (the $m_{s}=-1$ frequency and, for some spins, a local shift $\Delta_{i}$), which provides information for $A_{\parallel}$ based on the local shift ($\Delta_{i}\approx A_{\parallel}$), but not for $A_{\perp}$. For more information on the experimental procedures for spins 28-50, see [35], and see the caption of Table LABEL:tab:hf_compare for more details on which approximations were used to compute hyperfine couplings for specific spins. This variation in available experimental data highlights the practical motivation for inference methods that reduce the need for time-consuming full hyperfine characterization measurements.
@@
Some recovered spins have detection rates of zero (see Table LABEL:tab:hf_compare), meaning they are absent in all posterior samples. This occurs in the presence of sparse, noisy experimental data and inaccurate priors, highlighting the ill-posedness of the inverse problem: certain spin configurations cannot be reliably resolved given the limited information content. Importantly, the goal of this Bayesian inference approach is not exact, high-precision reconstruction but rather to provide informative posterior distributions that quantify uncertainty and identify plausible nuclear spin configurations. Hence, the method is particularly useful for rapid, high-throughput screening of spin environments, guiding experimental focus and helping prioritize spins for further investigation, rather than producing perfect reconstructions from limited data.
We do not provide explicit guidelines for a minimum detection rate required to consider a prediction “reliable,” because this threshold is highly context dependent. It can vary based on the specific experimental system, the quality and quantity of available data, and the application or target observable of interest. For example, in regimes with sparse or noisy data, even a moderate detection rate may provide useful information for screening or identifying likely spin candidates, whereas in high-precision applications, only spins with very high detection rates may be considered reliable. Similarly, the acceptable threshold may depend on whether one is interested in qualitative identification of spin presence, quantitative recovery of hyperfine parameters, or downstream tasks such as simulating spin-bath dynamics. In practice, the detection rate should be interpreted alongside other metrics, such as the baseline probability, posterior uncertainty, and consistency with prior knowledge, to assess the confidence in a given prediction.
Table 2: Comparison of Experimentally Measured and Recovered Hyperfine Interaction Values. This table compares experimentally measured and recovered hyperfine interaction values for nuclear spins, following the nuclear spin labeling convention in [16, 35]. The experimental data in this Table combines data from two different kinds of experimental measurements. The data for nuclear spins C1-C27 is from [16], in which the hyperfine components ($A_{\parallel,\text{exp}},A_{\perp,\text{exp}}$) are fit from direct measurements of the 13C spin precession frequencies for the $m_{s}=\pm 1$ electron spin projections (see Sec. I of the SI of [16] for more information about the experimental measurements and the fitting procedures). The data for nuclear spins C28-C50 is from [35], in which only the spin frequencies $(A^{-}_{\text{exp}})$ corresponding to the $m_{s}=-1$ electron projection are measured, and for a subset of these nuclear spins (denoted by ), a hyperfine shift $\Delta_{i}$ is measured (see Supplementary Note I of [35] for more details on experimental measurements and a derivation of the following approximations). For the nuclear spins denoted by , we approximate $A_{\parallel,\text{exp}}\approx\Delta_{i}$, and we recover $A_{\perp,\text{exp}}$ based on $A^{-}_{\text{exp}}=\sqrt{(\omega_{L}-A_{\parallel,\text{exp}})^{2}+A_{\perp,\text{exp}}^{2}}$, and any imaginary numbers are reported as 0. For the spins (denoted with ) where $\Delta_{i}$ was not measured (due to low polarization, coherence, or readout contrast), we approximate $A_{\parallel,\text{exp}}\approx\omega_{L}-A^{-}_{\text{exp}}$, where we have assumed that $A_{\parallel,\text{exp}}$ is small and $A_{\perp,\text{exp}}\approx 0$. These experimental values are compared with the hyperfine components ($A_{\parallel,\text{rec}},A_{\perp,\text{rec}}$) recovered using a hybrid MCMC algorithm (see Sec. 5.3). We report $A^{-}_{\text{rec}}$ as $A^{-}_{\text{rec}}=\sqrt{(\omega_{L}-A_{\parallel,\text{rec}})^{2}+A_{\perp,\text{rec}}^{2}}$. Detection rates of individual nuclear spins are determined based on the frequency of their appearance in spin configurations sampled from the posterior distribution. Spins can only be identified up to a degree of symmetry determined by the number of lattice sites with identical hyperfine couplings; however, we properly account for the multiplicity of spins with identical hyperfine couplings but different lattice sites when computing detection rates. Spins with detection rates above their baseline probabilities are marked with an asterisk (*).